Diazepam is used as a skeletal muscle relaxant and as a sedative administered before surgery, as well as a treatment for anxiety. For children above the age of 6 months who suffer from anxiety disorders, diazepam is administered at a reduced dosage. suppise a child with a body surface area of 0.8m^2 is to receive a dose of diazepam divided into three administrations. using the BSA rule and assuming that an adult dose of diazepam is 12 mg, what are the daily dose and the dose per administration?
Answer #1

BSA (Basal surface area) method of drug dose calculation is used two cancer patients and pediatric patients.

According to BSA ,

Patient’s dose = patient’s BSA (m2) X drug dose (mg) /1.73 m2

(1.73 m2 is average BSA of a adult)

So, child’s dose = 0.8 x 12 /1.73 = 5.54 or 5.5 mg

Hence, child’s daily dose of diazepam is 5.5 mg.

Dose per administration = total quality of dose (mg)/ number of doses

= 5.5/3 = 1.83 mg